An electric motor has a power rating of 1.60 kW and operates at 75% efficiency. The amount of work that it can do in 1.0 hrs is ___x10J. Give your answer with the correct number of sig digs and do not include units)


1
Expert's answer
2021-07-23T05:18:46-0400

By the definition of the power, we have:


"P_{total}=\dfrac{W}{t}."

From the definition of the efficiency, we have:


"\eta=\dfrac{P_{useful}}{P_{total}},""P_{useful}=\eta\cdot P_{total}."

Finally, we can find the amount of work done:


"W=\eta P_{total}t=0.75\cdot1.60\cdot10^{3}\ W\cdot3600\ s=4.32\cdot10^6\ J."